HETAV PATEL

Grade 11,

A 500g solid cube haivng an edge of length 20 cm floats in water.how much volume of the cube is outside the water?Density of water is 1000kg/m^3?

A 500g solid cube haivng an edge of length 20 cm floats in water.how much volume of the cube is outside the water?Density of water is 1000kg/m^3?

Grade:12

1 Answers

Nirmal Singh.
10 years ago
Suppose h be the height of cubical block above the water.
Volume of cubical block = 20 x 20 x 20 = 8x10^3 cm^3 = 8x10^3x 10^-6 m^3 = 8x10^-3 m^3
mass of cubical block = 500 g = 500x10^-3 = 0.5 kg
volume of water displaced by cubical block = 20x10^-2x20x10^-2x(20x10^-2 - h)
= 400x(20x10^-2-h)x10^-4 m^3
Mass of water displaced = 400x(20x10^-2 - h ) x10^-4 x 1000 = 40x(20x10^-2-h) kg
by law of floatation
Mass of water displaced = mass of cubical block
40x(20x10^-2-h) = 0.5
h = 0.075 m
h = 7.5 cm
Hence the volume above the water = 20x20x7.5 = 3000cm^3
Thanks & Regards,
Nirmal Singh
Askiitians faculty